OpenStudy (anonymous):

Sam scored 43, 48, and 42 points on three tests. How many points should he get on the fourth test to get an average score of 50? 83 points 67 points 45.75 points 44.33 points

OpenStudy (welshfella):

43 + 48 + 42 + x = 4*50 solve for x

OpenStudy (anonymous):

so the answer is..?

OpenStudy (welshfella):

If the average is 50 then the total marks for the 4 tests = 4*50 = 200

OpenStudy (welshfella):

i expect you can solve that equation